OpenStudy (anonymous):

A lamb grows rapidly, with its mass proportional to the cube of its length. When the lambs length changes by 15.8%, its mass increases by 17.3kg. Find the lambs mass at the end of the process.

OpenStudy (lalaly):

initial length=L mass=m m = kL^3 increase in length is 15.8% final length = L + 15.8L/100 = 1.158L final mass = m + 17.3 m + 17.3 = k x (1.158L)^3 m + 17.3 = 1.55kL^3 m + 17.3 = 1.55m 0.55m = 17.3 m = 31.45 final mass = 31.45 + 17.30 = 48.75 kg